How many integers between 15 and 75 are not evenly divisible by 3?
Answer by ikleyn(52946)

.

From 15 to 75 inclusive, there are  75 - 14 = 61 integer numbers.


There are 20 gaps of the length 3 between 15 and 75;  hence, the number of integers in this interval  [15,75] 
that are multiple of 3 is  20 + 1 = 21.


The numbers that are not divisible by 3 are the rest  61 - 21 = 40 numbers.    ANSWER

Solved.